OpenStudy (anonymous):

If E=4x-9,FG=3x-10, and EG=27, find the values of X, of EF and FG. PLEASE H E L P thanks :)

OpenStudy (kcmehne):

I do not believe this is solvable to a numerical answer. You have three equations and four unknowns. If you rearrange your equations with the following: 1a. EG=27 1b. G=27/E 2a. FG=3x-10 2b. F=(3x-10)/G Combine 1b and 2b F=(3x-10)/(27/E) -> F=(3x-10)E/27 insert E F=(3x-10)(4x-9)/27 therefore \[EF=(3x-10)(4x-9)^2/27\] Now your equations are the following: FG=3x-10 E=4x-9 \[EF=(3x-10)(4x-9)^2/27\] You still have three equations and four unknowns since finding X means finding or knowing E making your four unknowns X, FG, EF, and E. You need another constraining equation. Also this should probably go in the math group.